The Miami Central Hotel has 440 rooms. Every​ day, the housekeepers clean any room that was occupied the night before. If a guest is checking out of the​ hotel, the housekeepers give the room a thorough cleaning to get it ready for the next guest. This takes 25 minutes. If a guest is staying another​ night, the housekeeper only​ "refreshes" the​ room, which takes 20 minutes. Each​ day, each housekeeper reports for her 8​-hour ​shift, then prepares her cart. She pushes the cart to her floor and begins work. She usually has to restock the cart once per​ day; then she pushes it back to the storeroom at the end of the day and delivers dirty​ laundry, etc. Here is a​ timetable: ​1) Arrive at work and stock cart ​(0.10 ​hrs). ​2) Push cart to floor ​(0.10 ​hrs). ​3) Take morning break ​(0.33 ​hrs). ​4) Stop for lunch ​(0.50 ​hrs). ​ 5) Restock cart ​(0.30 ​hrs). ​6) Take afternoon break ​(0.33 ​hrs). ​7) Push cart back to laundry and store items ​(0.33 ​hrs). Last​ night, the hotel was full​ (all 440 rooms were​ occupied). People are checking out of 220 rooms. Their rooms will need to be thoroughly cleaned. The other 220 rooms will need to be refreshed. ​a) The number of minutes per day that each housekeeper can perform actual room cleaning is nothing minutes ​(enter your response as a whole​ number).
Answer #1

Total time per housekeeper = 480 minutes ( 8 hours)

Total time spent in Restocking + Break = 0.10+0.10+0.33+0.50+0.30+0.33+0.33 = 1.99 hours = ~2 hours or 120 minutes

number of minutes per day that each housekeeper can perform actual room cleaning is = 480 - 120 = 360 minutes